OpenStudy (anonymous):

A country's population in 1992 was 50 million. In 2002 it was 53 million. Estimate the population in 2010 using the exponential growth formula. Round your answer to the nearest million.

OpenStudy (anonymous):

can u look at mine real quick

OpenStudy (mathmagician):

\[53=50e^(10k)\], 10k=ln(53/50)=0.058, k=0.0058\[N(2010)=50*e^(0.0058*18)=55.5\] milions